OpenStudy (anonymous):

Arrange the following fractions in order from least to greatest. 1/3 , 2/5 , 3/8 A. 1/3 , 2/5 , 3/8 B. 1/3 , 3/8 , 2/5 C. 2/5 , 3/8 , 1/3 D. 3/8 , 1/3 , 2/5 [I think B]

OpenStudy (anonymous):

Between which two consecutive integers does\[\sqrt{57}\]lie? A. 7 and 8 B. 6 and 7 C. 9 and 10 D. 8 and 9 [I think its A]

OpenStudy (anonymous):

Arrange the following fractions in order from least to greatest. 1/3 , 2/5 , 3/8 A. 1/3 , 2/5 , 3/8 B. 1/3 , 3/8 , 2/5 C. 2/5 , 3/8 , 1/3 D. 3/8 , 1/3 , 2/5 this is B you are right

OpenStudy (anonymous):

Yay :)

OpenStudy (anonymous):

The links in your question wont load sorry i cant help you with that

OpenStudy (anonymous):

oh ok its cool um...

OpenStudy (anonymous):

Simplify completely: \[\sqrt{22y^8}\] A.\[2y^4\sqrt{5}\] B.\[10y^4\] C.\[2y^3\sqrt{5y}\] D.\[4y^4\sqrt{5}\]

OpenStudy (anonymous):

B is your answer for the first one! Message me:)

OpenStudy (anonymous):

i haven't done this yet sorry X{

OpenStudy (anonymous):

aw its cool dont worry :) you tired ;)

OpenStudy (anonymous):

Thanks for understanding and omg how did you know are you a wizard lol XD

OpenStudy (anonymous):

I guessed XD LIKE A BOSS! XD hahaa

OpenStudy (anonymous):

and like 47% wizards but shhh no one should ever know

OpenStudy (anonymous):

hahaha lol your seqrets safe with me P.s im 76% Pegasus but no one can know that either.....lol

OpenStudy (anonymous):

ahahahah XD you fun! lol

OpenStudy (anonymous):

thanks your fun to lol XP